• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 04:29:38 on April 19, 2011

    Entry number 356293

    keyword=Summary of R-HRS Optics Data taking during this owl shift

    We took some optics data on R-HRS while taking Produciton Data on L-HRS.
    
    E0=3.356, P0_R=2.600 (RQ3=2.2808),Theta=23.00
    
    1, Momentum Scan: Raster On, 4cm Dummy
    #3695: +3%: P_R=2.678,RQ3=2.3492, 20uA, 500K events  
    #3698: +0%: P_R=2.600,RQ3=2.2808, 20uA, 576K events 
    #3704: -3%: P_R=2.678,RQ3=2.3492, 10uA, 426K events
    
    2, Vertex Data: P0_R=2.600,RQ3=2.2808,Theta=23.00
    #3699, BeO target, Raster Off, 20uA, 400K
    #3700, Multi-C target, Raster Off, 20uA, 400K
    #3701, Multi-C target, Raster On, 20uA, 400K
    
    From Fig1,2, you can see the under our kinematics setting, the R-HRS
    optics matrix that we are using is worse than previous kinemtics setting
    during d-Threshold:
      http://www.jlab.org/~adaq/halog/html/1104_archive/110412163131.html
    
    Hence we need to get good optics data to calibrate the optics to get
    better resolution. So far the only data set we miss is Sieve Slit data. 
    


    A copy of this log entry has been emailed to: solvigno, doug, johna@anl.gov



    Run#3699 for BeO



    Run#3700 for Multi-C